The salt cesium chloride is soluble in water. When 11.9 g CsCl is dissolved in 117.00 g water, the temperature of the solution decreases from 25.00 °C to 22.56 °C. Based on this observation, calculate the dissolution enthalpy, ΔdissH, of CsCl. Assume that the specific heat capacity of the solution is 4.184 J g-1 °C-1 and that the energy transfer to the calorimeter is negligible.

ΔdissH = kJ/mol
Answer #1

Total mass of the sample = 11.9+117

= 128.9 grams

Q = mass * specific heat*∆T

= 128.9 * 4.184 * ( 25-22.6)

= 1294.36 joules

Moles of CsCl = mass / molar mass

= 11.9 / 168

= 0.071

Now, 0.071 mole absorbs 1294.36 joules

1 mole will absorb = 1294.36 / 0.071

= 18230.423 joules

Or 18.230 KJ/mol

If you have any query please comment

If you satisfied with the solution please like it thankxx